# Robust Matrix Completion for Discrete Rating-Scale Data

### v1

@@ 2.3 Column centering and initialization
Since the algorithm contains a soft-thresholded SVD step, it is important that the incomplete data matrix $\bm{X}$ is centered. One possibility is to center each column by the midpoint of the rating scale (i.e., the mean of the minimum and maximum rating category). While this can be expected to work well if the rating distributions of the columns are symmetric around the midpoint, such a setting is unrealistic in practical applications. In recommender systems applications, for instance, the distributions of popular items are typically skewed towards higher ratings, with the maximum rating often being the most frequent. In survey-based research, skewness towards one side of the rating scale is common, for instance, in organizational and management research (cf. Becker et al.,, 2019; Alfons et al.,, 2022). Hence, we center each column in the given data matrix by the median of its observed cells, and transform the set of rating categories $C_{j}$ accordingly for $j=1,\dots,p$, highlighting the need for a procedure that allows for different rating categories in different columns.
Using a median-centered matrix $\bm{X}$, we initialize the matrix $\bm{L}=P_{\Omega}(\bm{X})$ (corresponding to median imputation) and the discrepancy parameter $\bm{\Theta}$ as an $(n\times p)$-dimensional matrix of zeros. In practice, however, the algorithm is typically applied for a grid of values for $\lambda$ (see Section 2.5). Then the obtained solutions for $\bm{L}$ and $\bm{\Theta}$ are used as starting values for the next value of $\lambda$. As $\bm{\Theta}$ adjusts for the discrepancy between $\bm{L}$ and $\bm{Z}$ (which should increase with increasing $\lambda$), the values of $\lambda$ should thereby be sorted in ascending order.
@@ 2.4 Loss functions
For simplicity, we assume that the range of the rating categories is equal for all columns of $\bm{X}$, i.e., we have $c_{K}^{j}-c_{1}^{j}=r$ for $j=1,\dots,p$. In general, this can easily be enforced by normalizing the range before centering the columns, ensuring that all columns contribute similarly to the objective function in (2). We consider the following robust loss functions:
  • The pseudo-Huber loss $\rho(y)=\tau^{2}(\sqrt{1+(y/\tau)^{2}}-1)$. The inclusion of this loss function is motivated by its successful application in autoencoder neural networks for the detection of careless responding in rating-scale survey data (Alfons & Welz,, 2024; Welz & Alfons,, 2024). Here, we suggest to link the parameter $\tau$ to the average step size between rating categories by setting $\tau=r/(K-1)$.
@@
  • A bounded variant of the absolute loss given by $\rho(y)=\min(|y|,\tau)$ with $\tau=r/2$. That is, the loss is bounded at half the range of the rating categories.
Note that since the update step for $\bm{L}$ yields separable problems for its elements (see Section 2.2), the use of a nonconvex function such as the bounded absolute loss comes at no cost to computational complexity. A nonconvex loss may yield greater robustness in the selection of the regularization parameter $\lambda$ when minimizing the loss on a test set, which is discussed in the next section.
@@ 2.5 Selection of the regularization parameter
In order to select the regularization parameter $\lambda$ from a grid of candidate values based on out-of-sample prediction performance, some of the observed elements of $\bm{X}$ can be set to missing values for subsequent use as a validation set. For measuring the prediction error on the validation set, a natural choice is to apply the same loss function $\rho$ that is used for fitting the algorithm on the training set. It is possible to apply a cross-validation scheme whereby the observed elements are randomly divided into blocks, with each block being used as validation set once. However, even if the rows are independent, elements within the same row are not. It is therefore unclear if cross-validation would reduce the correlations among prediction errors on the different validation sets, compared to repeated holdout validation whereby a proportion of the observed elements are randomly selected in each replication to form the validation set. Hence, we prefer repeated holdout validation, as the proportion of observations in the validation set and the number of replications can be chosen independently.
@@ 3 Simulations
We now turn to evaluating the proposed procedure via simulations. We first present the methods included in the comparison, followed by two different simulation designs: one motivated by recommender systems, and one motivated by surveys in the social and behavioral sciences. Each setting uses 100 replications.
3.1 Methods
For the proposed method (denoted by RDMC for robust discrete matrix completion), we consider the three loss functions described in Section 2.4 (denoted by pHuber for the pseudo-Huber loss, absolute for the absolute loss, and bounded for the bounded absolute loss) and set the parameters of the algorithm as described in Section 2.2. We compare RDMC with Soft-Impute (SI) (Mazumder et al.,, 2010; Hastie et al.,, 2015), as well as with a variant that discretizes the obtained predictions to the given rating scale via the mapping function $m_{j}(y)=\min(\max([y],c_{\min}),c_{\max})$, where $c_{\min}$ and $c_{\max}$ are the minimum and maximum rating, respectively, and []delimited-[][$\cdot$][ ⋅ ] denotes rounding to the nearest integer (SI-discretized). We use the SVD-based algorithm due to its theoretical convergence guarantees and set the same convergence threshold as for RDMC. For selecting the regularization parameter $\lambda$, both RDMC and SI use repeated holdout validation with 5 replications and 10% of the observed cells to be randomly selected as validation set. The candidate values are thereby given by a logarithmic grid of ten values between 0.01 and 1, which are scaled by the largest singular value of $P_{\Omega}(\bm{X}_{\text{train}})$, with $\bm{X}_{\text{train}}$ denoting the median-centered (for RDMC) or mean-centered (for SI) training data. Additional benchmark methods are median imputation (denoted by median), a discretized variant of median imputation (median-discretized; in case the median of a column falls in between rating categories, the predictions are randomly sampled from the corresponding two categories), and mode imputation (mode; in case of a column with multiple modes, one of them is selected at random for each missing cell).
3.2 Simulations motivated by recommender systems
3.2.1 Data Generation
We simulate data of $n=300$ users rating $p=200$ items. We start by simulating latent continuous data from the low-rank matrix factorization model $\bm{Z}=\bm{A}\bm{B}^{\top}+\mathcal{E}$, where $\bm{A}$ is of dimension $n\times q$, $\bm{B}$ of dimension $p\times q$, and $\bm{\mathcal{E}}$ of dimension $n\times p$, with rank $q=20$ and the elements of $\bm{A}$, $\bm{B}$, and $\bm{\mathcal{E}}$ being independent and standard normally distributed. Subsequently, we rescale $\bm{Z}^{*}=\bm{Z}/\sqrt{q+1}$ so that the elements of $\bm{Z}^{*}$ have variance 1.
@@ 3.2.1 Data Generation
For discretizing the resulting continuous data matrix in order to obtain a rating matrix $\bm{X}$, we use the following breakpoints depending on the number of rating categories:
@@
  • $K=5$: to simulate common 1 to 5 star ratings, we set the breakpoints at $-1.5$, $-0.5$, $0.5$, and $1.5$.
  • $K=10$: to simulate common 1 to 10 star ratings (or, equivalently, ratings up to 5 stars but allowing for half-stars), we set the breakpoints at $-2$, $-1.5$, $-1$, $-0.5$, 00, $0.5$, $1$, $1.5$, and $2$.
3.2.2 Missing values
We generate missing values in the rating matrix $\bm{X}$ using two different settings:
  • Missing not at random (MNAR): The negated mean shifts $-s_{1},\dots,-s_{p}$ are mapped to the interval $[0.4,0.99]$ in order to obtain the proportion of observations in each item to be replaced by missing values. For instance, the most popular item (large positive mean shift, high ratings) contains 40% missing values, while the most unpopular item (large negative mean shift, low ratings) contains 99% missing values. [^1] These choices are motivated by the MovieLens 100K data (Harper & Konstan,, 2015) used in Section 4.1, in which the most complete item contains $38.2\%$ missing values and the most incomplete item contains more than 99% missing values. Hence, low ratings are much more likely to be missing than high ratings.
  • Missing completely at random (MCAR): a proportion $\gamma=0.7$ of cells in $\bm{X}$ is randomly selected and replaced by missing values. Note that $\gamma$ is chosen so that the overall proportion of missing values is similar to the MNAR setting.
3.2.3 Attacks
To investigate the robustness of the methods against profile injection attacks (also known as shilling attacks), we focus on so-called nuke attacks aimed at demoting a certain target item, i.e., decreasing the probability of it being recommended. We apply three efficient attack schemes, denoted by average, reverse bandwagon, and love/hate (see Mobasher et al., (2007) for a detailed overview of common attack schemes).
@@ 3.2.3 Attacks
3.2.4 Evaluation criteria
@@
$MAE=\frac{1}{|\Gamma|}\sum_{(i,j)\in\Gamma}|X_{ij}-\hat{X}_{ij}|,$ (3)
where $\Gamma$ denotes the index set of missing cells in $\bm{X}$, and $\hat{X}_{ij}$ is is the prediction of the cell $X_{ij}$. For the different attack settings, we evaluate the methods via the mean prediction shift (MPS) in the target variable (e.g., Mobasher et al.,, 2007), i.e.,
$MPS=\frac{1}{|\Gamma_{\text{target}}|}\sum_{i\in\Gamma_{\text{target}}}\hat{X}_{i,j_{\text{target}}}^{\text{after}}-\hat{X}_{i,j_{\text{target}}}^{\text{before}},$
where $\Gamma_{\text{target}}$ is the index set of missing cells in the target variable $j_{\text{target}}$ of $\bm{X}$, while $\hat{X}_{ij}^{\text{before}}$ and $\hat{X}_{ij}^{\text{after}}$ denote the prediction of the cell $X_{ij}$ before and after the attack, respectively.
3.2.5 Results
For RDMC, we focus on the pseudo-Huber loss, as preliminary simulations indicate that results for the absolute loss and the bounded absolute loss are similar (see Appendix A.1). Figure 1 displays box plots of the MAE for the setting without an attack. In case of three rating categories, SI without the discretization step does not improve upon median and mode imputation. The discretized variant, on the other hand, yields a clear improvement in MAE, with RDMC falling in between. For five rating categories, SI and RDMC considerably outperform the benchmark methods, with RDMC being in between the two SI approaches for the MNAR setting, but performing equally well as SI-discretized in the MCAR setting. For ten rating categories, RDMC yields the best performance in terms of MAE, while both SI approaches exhibit large variability in the MCAR setting.
@@ 3.2.5 Results
We now turn to stability of the predictions under attacks, with Figure 2 containing box plots of the MPS for the MNAR setting. Results for the MCAR setting are qualitatively similar and can be found in Appendix A.1. Both SI approaches are strongly influenced by the attacks, with the average attack having the biggest impact. The MPS is thereby large enough that the target item would be recommended to far fewer users—for instance, the predictions on average drop by at least one full rating category in the setting with ten categories. RDMC is far more stable in the presence of attacks, with an MPS much closer to zero in all settings. Furthermore, while median regression is stable for three and five rating categories, it yields an MPS of -1 in most instances with ten rating categories.
@@
On a final note, we also included a variant of RDMC with the pseudo-Huber loss that we do not iterate until convergence, but for which we perform at most 10 iterations (denoted by RDMC-pHuber-10 in the figures). Interestingly, the MAE in case of no attack is often smaller compared to iterating until convergence, whereas the MPS in the presence of an attack is slightly larger in absolute value in case of three rating categories.
3.3 Simulations of survey data
3.3.1 Data generation
We simulate rating-scale responses of $n=300$ participants to a survey on $q=10$ latent constructs, with each construct being measured by $r\in\{4,8\}$ items such that the total number of items is $p\in\{40,80\}$. For this purpose, we first simulate the underlying latent sentiments $\bm{Z}$ from a multivariate normal distribution $N(0,\bm{\Sigma})$, where $\bm{\Sigma}$ follows a block-Toeplitz structure with 1 on the diagonal and the remaining elements being given by $\rho_{kl}=0.6^{|k-l|+1}$ with $k=1,\dots,q$ and $l=1,\dots,q$ denoting the row and column indices of the blocks in $\bm{\Sigma}$.
To generate items with skewed distributions, we add random mean shifts to the columns of $\bm{Z}$—with all items within the same construct receiving the same mean shift—before discretization into $K\in\{5,7,9\}$ answer categories interpreted as values $\{1,\dots,K\}$. Similar to the previous simulations, the mean shifts are drawn from the interval $[0,s_{\text{max}}]$, where $s_{\text{max}}$ is chosen to result in 40% of the responses being expected in the highest answer category.
The resulting continuous data matrix is then discretized using equispaced breakpoints $-K/2+1,-K/2+2,\dots,K/2-2,K/2-1$. To simulate reverse-keyed items, which are commonly used in practice, half the items of each construct are reversed by reassigning answer category 1 to $K$, 2 to $K-1$, and so on. Note that this implies that the data contains both left-skewed and right-skewed items. As is common in the behavioral sciences, we randomize the order of the items in the survey (with the same order being used for all participants), resulting in the final rating-scale data matrix $\bm{X}$.
3.3.2 Missing values and careless responding
We inject missing values into the data matrix $\bm{X}$ by simulating respondents who stop responding at some point in the survey. We set the proportion of respondents who abandon the survey to $\gamma\in\{0.2,0.4,0.6\}$. For each of those respondents, we randomly select the item at which they stop responding, with all responses from this item onwards being replaced by missing values. We emphasize that here $\gamma$ does not indicate the proportion of missing cells in $\bm{X}$, but the number of rows that contain missing values.
To simulate careless respondents, a proportion $\varepsilon\in\{0,0.1,0.2\}$ of the rows in $\bm{X}$ are replaced with careless respondents who randomly select from the two extreme answer categories $\{1,l\}$. Note that our simulation allows for the situation that a careless respondent abandons the survey such that a certain row may contain both careless responses and missing values.
3.3.3 Evaluation criteria
3.3.4 Results
We focus on the results for the settings with 20% of respondents who abandon the survey, as the results for 40% and 60% are similar (see Appendix A.2). Box plots for the MAE in the simulated survey with $p=40$ items are shown in Figure 3. In general, differences between the methods are smaller than in the simulations for recommender systems. On the other hand, we now see more differences among RDMC with different loss functions. Most notably, RDMC with the pseudo-Huber function remains close to SI-discretized in the absence of careless responding, while RDMC with the bounded absolute loss yields the best results for 20% careless respondents. In addition, with a higher number of answer categories, robustness benefits of RDMC over SI become more apparent. For instance, with 10% of careless respondents, RDMC and SI-discretized perform similarly for five answer categories, whereas all three variants of RDMC outperform SI-discretized in case of seven or nine answer categories.
Figure 3: Results for the simulated survey with $p=40$ variables and 20% of respondents who abandon the survey. The rows correspond to different number of answer categories and the columns to different proportions of careless respondents.
Increasing the number of items to $p=80$, we now observe bigger differences between the methods. Figure 4 contains the corresponding box plots of the MAE. For all settings regarding the number of answer categories, the MAE of RDMC remains in between that of the two SI approaches in case of no careless responding, while RDMC yields better performance in the presence of careless respondents. As previously, the pseudo-Huber loss is preferable in the absence of careless responding (with an MAE close to that of SI-discretized), while the bounded absolute loss is the most robust for larger prevalence of careless respondents.
Figure 4: Results for the simulated survey with $p=80$ variables and 20% of respondents who abandon the survey. The rows correspond to different number of answer categories and the columns to different proportions of careless respondents.
Finally, we included variants of RDMC with the three loss functions for which we perform at most 10 iterations. Here, the MAE is at best similar but in most cases slightly higher compared to iterating until convergence. For simplicity, we thus omit these variants from Figures 3 and 4 (as well as Figures A.4A.7 in Appendix A.2).
4 Empirical Applications

### v2

@@ 2.3 Column centering and initialization
Since the algorithm contains a soft-thresholded SVD step, it is important that the observed (incomplete) rating matrix $\bm{R}$ is centered. One possibility is to center each column by the midpoint of the rating scale (i.e., the mean of the minimum and maximum rating category). While this can be expected to work well if the rating distributions of the columns are symmetric around the midpoint, such a setting is unrealistic in practical applications. In recommender systems, the distributions of popular items, for instance, are typically skewed towards higher ratings, with the maximum rating often being the most frequent. Hence, we center the $j$th column in the given data matrix $\bm{R}$ by the median $M_{j}$ of its observed cells, i.e., we obtain $\bm{X}$ by setting $X_{ij}=R_{ij}-M_{j}$, $i=1,\dots,n$, $j=1,\dots,p$. Accordingly, we transform the set of rating categories to $C_{j}=\{c_{1}^{j},\dots,c_{K}^{j}\}$ with $c_{k}^{j}=k-M_{j}$ for $k=1,\dots,K$. This highlights the need for a procedure that allows for different rating categories in different columns, a feature that our proposed procedure RDMC accommodates.
Using the median-centered matrix $\bm{X}$, we initialize the matrix $\bm{L}=P_{\Omega}(\bm{X})$ (corresponding to median imputation) and the discrepancy parameter $\bm{\Theta}$ as an $(n\times p)$-dimensional matrix of zeros. However, the algorithm is typically applied for a grid of values for $\lambda$ (see Section 2.5). Then the obtained solutions for $\bm{L}$ and $\bm{\Theta}$ for a given value of $\lambda$ are used as starting values for the next value of $\lambda$. As $\bm{\Theta}$ adjusts for the discrepancy between $\bm{L}$ and $\bm{Z}$ (which should increase with increasing $\lambda$), the values of $\lambda$ should thereby be sorted in ascending order.
@@ 2.4 Loss functions
In order to reduce the influence of large errors due to corrupted observations such as fake profiles, we consider the following robust loss functions for the (pseudo-)norm $\|\cdot\|_{\rho}$ in (2), which are common choices in the literature on robust methods:
  • The pseudo-Huber loss $\rho(y)=\tau^{2}(\sqrt{1+(y/\tau)^{2}}-1)$ with parameter $\tau$. The inclusion of this loss function is motivated by its successful application to rating-scale data in a different context, namely in autoencoder neural networks for the detection of careless responding in surveys from the behavioral sciences (Alfons & Welz,, 2024; Welz & Alfons,, 2025). Here, we suggest to link the parameter $\tau$ to the step size between rating categories by setting $\tau=1$.
@@
  • A truncated variant of the absolute loss given by $\rho(y)=\min(|y|,\tau)$. We suggest to set $\tau=(K-1)/2$, i.e., the loss is truncated at half the range of the rating categories.
Since the update step for $\bm{L}$ yields separable problems for its elements (see Section 2.2), the use of a nonconvex function such as the truncated absolute loss comes at no cost to computational complexity. A nonconvex loss may yield greater robustness in selecting the regularization parameter $\lambda$ when minimizing the loss on a test set, which is discussed next.
@@ 2.5 Selection of the regularization parameter
To select the regularization parameter $\lambda$ from a grid of candidate values based on out-of-sample prediction performance, some of the observed elements of $\bm{R}$ can be set to missing values for subsequent use as a validation set. For measuring the prediction error on the validation set, a natural choice is to apply the same loss function $\rho$ that is used for fitting the algorithm on the training set. It is possible to apply a cross-validation scheme whereby the observed elements are randomly divided into blocks, with each block being used as validation set once. However, even if the rows are independent, elements within the same row are not. It is therefore unclear if cross-validation would reduce the correlations among prediction errors on the different validation sets, compared to repeated holdout validation whereby a proportion of the observed elements are randomly selected in each replication to form the validation set. Hence, we prefer repeated holdout validation, as the proportion of observations in the validation set and the number of replications can be chosen independently.
@@ 3 Simulations
Before applying the proposed procedure in empirical case studies, we thoroughly evaluate it via simulations. Crucially, our simulation design closely mimics recommender systems applications so that it may serve as blueprint for future studies. We perform 100 replications.
3.1 Data generation
We simulate data of $n=300$ user ratings on $p=200$ items. We start by simulating latent continuous data from the low-rank matrix factorization model $\bm{Z}=\bm{A}\bm{B}^{\top}+\mathcal{E}$, where $\bm{A}$ is of dimension $n\times q$, $\bm{B}$ of dimension $p\times q$, and $\bm{\mathcal{E}}$ of dimension $n\times p$, with rank $q=20$ and the elements of $\bm{A}$, $\bm{B}$, and $\bm{\mathcal{E}}$ being independent and standard normally distributed. Subsequently, we rescale $\bm{Z}^{*}=\bm{Z}/\sqrt{q+1}$ so that the elements of $\bm{Z}^{*}$ have variance 1.
@@ 3.1 Data generation
For discretizing the resulting continuous data matrix in order to obtain a rating matrix $\bm{R}$, we use the following breakpoints depending on the number of rating categories:
@@
  • $K=5$: to simulate common 1 to 5 star ratings, we set the breakpoints at 1.5-1.5- 1.5, 0.5-0.5- 0.5, 0.50.50.5, and 1.51.51.5.
  • $K=10$: to simulate common 1 to 10 star ratings (or, equivalently, ratings up to 5 stars but allowing for half-stars), we set the breakpoints at 2-2- 2, 1.5-1.5- 1.5, 1-1- 1, 0.5-0.5- 0.5, 0, 0.50.50.5, 111, 1.51.51.5, and 222.
3.2 Missing values
We generate missing values in the rating matrix $\bm{R}$ using two different settings:
[…]
  • Missing completely at random (MCAR): a proportion $\gamma=0.7$ of cells in $\bm{R}$ is randomly selected and replaced by missing values. Here, $\gamma$ is chosen so that the overall proportion of missing values is similar to the MNAR setting. Note that this setting is included for reference purposes since MCAR is unrealistic in recommender systems applications.
3.3 Attacks
To investigate the robustness of the methods against adversarial manipulation with profile injection attacks (also known as shilling attacks), we focus on so-called nuke attacks aimed at demoting a certain target item, i.e., decreasing the probability of it being recommended. We apply three efficient attack schemes, denoted by average, reverse bandwagon, and love/hate (see Mobasher et al.,, 2007 for a detailed overview of common attack schemes).
@@ 3.3 Attacks
3.4 Methods
For the proposed method RDMC, we consider the three loss functions described in Section 2.4 (denoted by pHuber for the pseudo-Huber loss, absolute for the absolute loss, and truncated for the truncated absolute loss) and set other parameters as described in Section 2.2.
We compare RDMC with the popular procedure Soft-Impute (SI) (Mazumder et al.,, 2010; Hastie et al.,, 2015). Since this method yields continuous predictions, we introduce a variant that discretizes the obtained predictions to the given rating scale via the mapping function $m_{j}(y)=\min(\max([y],c_{\min}),c_{\max})$, where $c_{\min}$ and $c_{\max}$ are the minimum and maximum rating, respectively, and [][$\cdot$][ ⋅ ] denotes rounding to the nearest integer (SI-discretized). We use the SVD-based algorithm due to its theoretical convergence guarantees and set the same convergence threshold as for RDMC.
For selecting the regularization parameter $\lambda$, both RDMC and SI use repeated holdout validation with 5 replications and 10% of the observed cells to be randomly selected as validation set. The candidate values are thereby given by a logarithmic grid of ten values between 0.01 and 1, which are scaled by the largest singular value of $P_{\Omega}(\bm{X}_{\text{train}})$, with $\bm{X}_{\text{train}}$ denoting the median-centered (for RDMC) or mean-centered (for SI) training data.
Additional benchmark methods are median imputation (denoted by median), a discretized variant of median imputation (median-discretized; in case the median of a column falls in between rating categories, the predictions are randomly sampled from the corresponding two categories), and mode imputation (mode; in case of a column with multiple modes, one of them is selected at random for each missing cell).
3.5 Evaluation criteria
@@
$MAE=\frac{1}{|\Gamma|}\sum_{(i,j)\in\Gamma}|R_{ij}-\widehat{R}_{ij}|,$ (5)
where $\Gamma$ denotes the index set of missing cells in $\bm{R}$, and $\widehat{R}_{ij}$ is the prediction of the cell $R_{ij}$. For the different attack settings, we evaluate the methods via the mean prediction shift (MPS) in the target variable (e.g., Mobasher et al.,, 2007), i.e.,
$MPS=\frac{1}{|\Gamma_{\text{target}}|}\sum_{i\in\Gamma_{\text{target}}}\left(\widehat{R}_{ij_{\text{target}}}^{\text{after}}-\widehat{R}_{ij_{\text{target}}}^{\text{before}}\right),$
where $\Gamma_{\text{target}}$ is the index set of missing cells in the target variable $j_{\text{target}}$ of $\bm{R}$, while $\widehat{R}_{ij}^{\text{before}}$ and $\widehat{R}_{ij}^{\text{after}}$ denote the prediction of the cell $R_{ij}$ before and after the attack, respectively.
3.6 Results
@@
For RDMC, we focus on the pseudo-Huber loss, as preliminary simulations indicate that results for the absolute loss and the truncated absolute loss are similar (see Appendix A). Figure 1 displays box plots of the MAE for the setting without an attack. In case of three rating categories, SI without the discretization step does not improve upon median and mode imputation. The discretized variant, on the other hand, yields a clear improvement in MAE, with RDMC falling in between. For five rating categories, SI and RDMC considerably outperform the benchmark methods, with RDMC being in between the two SI approaches for the MNAR setting, but performing equally well as SI-discretized in the MCAR setting. For ten rating categories, RDMC yields the best performance in terms of MAE, while both SI approaches exhibit large variability in the MCAR setting.
@@
We now turn to stability of the predictions under adversarial manipulation, with Figure 2 containing box plots of the MPS for the MNAR setting. Results for the MCAR setting are qualitatively similar and can be found in Appendix A. Both SI approaches are strongly influenced by the attacks, with the average attack having the biggest impact. The MPS is thereby large enough that the target item would be recommended to far fewer users—for instance, the predictions on average drop by at least one full rating category in the setting with ten categories. RDMC is far more stable in the presence of attacks, with an MPS much closer to zero in all settings. Furthermore, while median regression is stable for three and five rating categories, it yields an MPS of 1-1- 1 in most instances with ten rating categories.
@@
On a final note, we also included a variant of RDMC with the pseudo-Huber loss that we do not iterate until convergence. Instead, we apply a liberal stopping criterion, limiting the procedure to a maximum of 10 iterations (denoted by RDMC-pHuber-liberal in the figures). Interestingly, the MAE in case of no attack is often smaller compared to iterating until convergence, whereas the MPS in the presence of an attack is slightly larger in absolute value in case of three rating categories.
4 Empirical case studies